ghostmasterTy:

1. 4/16 = 2. 9/27 = 3. 7/49 = 4. 8/64 = 5. 3/21 = Simplify these improper fractions 6. 45/90 = 7. 203/4 = 8. 87/4 = Multiply these mixed numbers. You have to make the mixed numbers into improper fractions first. You do this by multiplying the denominator by the whole number, and then add the numerator. 9. 2 ⅞ x 3 ⅘ = 10. 1 4/5 x 2 1/9 = 11. 1 5/6 x 2 1/3 = you dont have math

snowflake0531:

Hold on hold on, let's first look at the first 5

ghostmasterTy:

ok

snowflake0531:

1. 4/16 = \[\frac{ 4 }{ 16 } \div \frac {4}{4}\]

ghostmasterTy:

and you dont have subject math

ghostmasterTy:

ok

snowflake0531:

There actually is a mathematics subject lol

ghostmasterTy:

ok

ghostmasterTy:

lets dob th math

snowflake0531:

@snowflake0531 wrote:
1. 4/16 = \[\frac{ 4 }{ 16 } \div \frac {4}{4}\]

ghostmasterTy:

4

snowflake0531:

-? What is 4/4, what is 16/4

ghostmasterTy:

i do not now

snowflake0531:

do you know what 4 divided by 4 is?

ghostmasterTy:

1 ]

ghostmasterTy:

im not dum

ghostmasterTy:

ok 4 d bbyn 4 os 1

ghostmasterTy:

ok

ghostmasterTy:

and

snowflake0531:

Okay so it simplifies to 1/4 yea?

ghostmasterTy:

3

snowflake0531:

huh-?

ghostmasterTy:

3

snowflake0531:

what is 3

ghostmasterTy:

1/3

snowflake0531:

Oh yes the second one is 1/3

ghostmasterTy:

1/4

snowflake0531:

the first one is 1/4 yes

ghostmasterTy:

so whats the answer

snowflake0531:

1/4

ghostmasterTy:

so where done with number 1

snowflake0531:

huh?

ghostmasterTy:

ok number 1 is 1/4

ghostmasterTy:

number to now

ghostmasterTy:

9/27

snowflake0531:

\[\frac{ 9 \div9 }{ 27 \div9 }\]

ghostmasterTy:

1/3

snowflake0531:

yes

snowflake0531:

\[\frac{ 7 \div 7 }{ 49 \div7 }\]

ghostmasterTy:

7/49

ghostmasterTy:

1/7

snowflake0531:

yes

snowflake0531:

\[\frac{ 8 \div 8 }{ 64 \div 8 }\]

ghostmasterTy:

8/8

snowflake0531:

are you sure?

ghostmasterTy:

yes

snowflake0531:

what's 8 divided by 8

ghostmasterTy:

ok lets do this in private cuz i don't want anyone to think that im dum

ghostmasterTy:

oh

ghostmasterTy:

8/1

snowflake0531:

No 1/8

ghostmasterTy:

ok

snowflake0531:

\[\frac{ 3 \div 3 }{ 21 \div 3 }\]

ghostmasterTy:

7/1

ghostmasterTy:

im going to be youur fan

ghostmasterTy:

ok

snowflake0531:

It's actually the opposite, 1/7 and lol thx

ghostmasterTy:

so it 1/7

snowflake0531:

yes

ghostmasterTy:

oh i no the next one is it 1/9

ghostmasterTy:

no 1/5

snowflake0531:

no, number 6 is \[\frac{ 45 \div 45 }{ 49 \div 45 }\]

ghostmasterTy:

oh

ghostmasterTy:

ok 1/1

snowflake0531:

oops, i meant 90/45

ghostmasterTy:

ok

snowflake0531:

\[\frac{ 45 \div 45 }{ 90 \div 45 }\]

ghostmasterTy:

1/2

snowflake0531:

yes

snowflake0531:

\[\frac{ 203 }{ 4 } = \frac {4 \times 50 +3}{4} = 50 \frac {3}{4}\] do you see how i kinda got that-? lol

ghostmasterTy:

htis next one is hard

ghostmasterTy:

htis next one is hard

ghostmasterTy:

did you do it

snowflake0531:

@snowflake0531 wrote:
\[\frac{ 203 }{ 4 } = \frac {4 \times 50 +3}{4} = 50 \frac {3}{4}\] do you see how i kinda got that-? lol

ghostmasterTy:

yes

ghostmasterTy:

is it 3/4 or 4/3

snowflake0531:

Which one are you doing

ghostmasterTy:

203 /4

ghostmasterTy:

the end

snowflake0531:

you put \[50 \frac {3}{4}\]

ghostmasterTy:

ok

ghostmasterTy:

87/4

ghostmasterTy:

ok when you need help with times come to me ok

snowflake0531:

i gtg

ghostmasterTy:

you have help me and you know what you did you have tech me how to do this

snowflake0531:

o-o I'm confused by the grammar

ghostmasterTy:

ok

ghostmasterTy:

im done

snowflake0531:

you're done? Do you want me to check answers-

ghostmasterTy:

no